Circular No. 8312

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004az IN UGC 6853 Further to IAUC 8311, D. Singer and W. Li report the LOSS discov-

ery, on KAIT images taken on Mar. 21.32 and 23.39 UT, of an apparent supernova (mag ∼ 17.8) located at α = 11^{h}52^{m}48^{s}.22, δ = +29^{o}20′51′′.7 (equinox 2000.0), which is 11′′.9 east and 67′′.2 north of the nucleus of UGC 6853. A KAIT image taken on Jan. 29.40 showed nothing at this position (limiting mag ∼ 19.0).

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2004az, obtained by P. Berlind on Mar. 25.26 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova several weeks past maximum brightness.

SUPERNOVA 2004aw IN NGC 3997 S. Benetti, N. Elias-Rosa, G. Blanc, H. Navasardyan, M. Turatto, and

L. Zampieri, Osservatorio Astronomico di Padova; E. Cappellaro, Osservatorio Astronomico di Capodimonte; and M. Pedani, Telescopio Nazionale Galileo (TNG), on behalf of the ERTN (IAUC 7987), obtained a spectrum of SN 2004aw (cf. IAUC 8310, 8311) with the TNG (+ Dolores; range 335– 995 nm; resolution 1.4 nm) on Mar. 24.94 UT. The supernova has now evolved to resemble the spectrum of SN 1991T, a few days after maximum, and is therefore classified as type Ia. This also accounts for the fact that the narrow interstellar absorption line seen in the supernova spectrum indicates a reddening of E(B − V) ∼ 0.30 in the host galaxy. The expansion velocity deduced from the Si-II 635.5-nm feature is ∼ 12600 km/s.

SUPERNOVA 2004ay IN UGC 11255 Matheson et al. add that a spectrum of SN 2004ay (cf. IAUC 8311),

obtained by Berlind on Mar. 26.49 UT, shows it to be a type-IIn supernova.

SUPERNOVAE 2002lu AND 2002lv Further to IAUC 8311, T. Schrabback adds that the data analysis was

done with D. Wuttke, H. Knuth, and J.-M. Miralles.

2004 March 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green